### v17

# Rational angle bisectors on the coordinate plane and solutions of Pell’s equations

Abstract
On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author. In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers. We also describe every nontrivial rational solution of the above equation as one of its applications.
1 Introduction
On the coordinate plane, we consider the following problem, which can be called the rational angle bisection problem.
Problem 1.
For which rational numbers are the slopes of the angle bisectors between two straight lines with slopes rational?
Remark 1.
Given two straight lines, we consider the two angles formed by them, regardless of whether they are acute or not. The bisector of one of the angles and that of the supplementary angle are perpendicular to each other. In the case when they are not parallel to the coordinate axes, if one of the slopes is rational, then so is the other, since the product of the slopes is
Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques. Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without error due to approximations to irrational numbers by using a solution to this problem. Problem 1 is reduced to solving the equation
in (see [5, Proposition 1]). We say that a solution of ( ‣ 1) is trivial if
Example 1.
The following proposition plays an important role in solving Problem 1.
Proposition 1.
Every nontrivial rational solution of ( ‣ 1) is given by
for some rational solutions of
where is a square-free integer
The proof is given in Section 3. The following formula for integral solutions of ( ‣ 1) was proven by using certain criteria for the divisibility of integral solutions of Pell’s equations.
Theorem 1 ([5, Theorem 1]).
Throughout this article, we use the following definition and symbols.
Definition 1.
Let be a square-free integer and let be an integer
Let be an integral solution of We say that is strictly primitive if are coprime.
Let be positive integral solutions of We say is smaller than if
For each equation we call its minimum positive integral solution its fundamental solution.
Remark 2.
Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free. In particular, every integral solution of is strictly primitive.
Notation 1.
Let denote the additive monoid of all integers
Let be a square-free integer Let be the fundamental unit of the real quadratic field Let be the fundamental solution of and put
Put
For each put
and let be the fundamental solution of
and put
Put
Furthermore, for each where we denote the conjugate and norm of by
respectively.
The first main theorem of this article is the following formula.
Theorem 2.
Let be a nontrivial rational solution of ( ‣ 1).
Case : are the -components of rational solutions of the equation Then is given by
for some such that
Case : are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
for some such that and In addition, are written of the form
with some satisfying
and
The proof is given in Section 3.
Example 2.
The result in Case (I) refines the previous result [5, Theorem 2]. For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle. We find the rational solutions of ( ‣ 1) in Case (II) by using the following Theorem. This is the second main theorem of this article.
Theorem 3.
Let be an integer
Suppose that or “ and ” or Then has a strictly primitive integral solution if and only if
(1.5) for some for each prime number where
Suppose that and Then has a strictly primitive integral solution if and only if
(1.6) for some for each prime number
In these cases, its strictly primitive integral solution satisfies
(1.7) for some such that
(1.8) and where in the latter case.
- (2)
The proof is given in Section 2.
2 Rational solutions of Pell’s equations
In this section, we prove Theorem 3. Throughout the section, let be a square-free integer and use the symbols in Notation 1. We recall the following classical theorem.
Theorem 4.
The positive Pell’s equation has a nontrivial integral solution independently of the value of
The negative Pell’s equation has a rational solution, if and only if has no prime factors congruent to modulo
Proof.
See [1, Theorem 3.2.1].
By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime factors congruent to modulo We can assume since Under these conditions, is a rational solution of
Conversely, if has a rational solution with then which implies that has no prime factors congruent to modulo by Fermat’s theorem as above, and so does ∎
Remark 3.
If has an integral solution, then has no prime factors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
Let If have integral solutions respectively, then has an integral solution
since the norm map is a group homomorphism. In the case when such a solution is called a Pell multiple.
Theorem 5.
Let If has an integral solution, then there exists a finite number of its integral solutions such that its every integral solution satisfies
for some and
Proof.
See [4, Corollary 3.5]. ∎
In this article, we generalize the concept of Pell multiples. The following propositions are fundamental.
Proposition 2.
Let be distinct prime numbers, and let If has a strictly primitive integral solution for each then defined by
(2.1) is a strictly primitive integral solution of
Let be a prime number, and let be integers If has a strictly primitive integral solution then defined by
(2.2) is a strictly primitive integral solution of
Proof.
We prove the contrapositions.
For each let be an integral solution of and put Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
Case: Then and therefore
Case: Then This implies that divides
and therefore in Hence
Thus and therefore the solution is not strictly primitive.
Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then
Case: Then and therefore
Case: Then and therefore in With respect to an extension of the -adic additive valuation to this implies
Hence
Thus and therefore the solution is not strictly primitive.∎
Proposition 3.
Let Suppose that has an integral solution. If has a strictly primitive integral solution, then so does
Proof.
We prove the contraposition. Suppose that has no strictly primitive integral solutions. Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that
Case: Then and therefore
Case: Then In this implies and therefore since Hence
Thus This implies that has no strictly primitive integral solutions. ∎
We say that an ideal in or the integer ring of a quadratic field is primitive if is not divisible by for any prime number
Theorem 6.
Let be a prime number. Suppose that or Then has a strictly primitive integral solution for some multiple of if and only if splits in
Suppose that
If then
Both and hold if and only if In this case, holds.
If and is a multiple of then every integral solution of is a pair of even numbers, and not strictly primitive.
If or “ and ” or then
If and then
Unless “ and ” then
for each
Proof.
Put Recall that the integer ring of is
Let be the monoid of all ideals in and let be the equivalence relation on defined by
which is compatible with the multiplication. Since the ideal class group of is finite, and so is the quotient monoid which can be regarded as its submonoid. Furthermore, splits in if and is unramified in if
Suppose that a prime number splits into prime ideals and in that is, Then there exist such that
by the finiteness of where are coprime since is primitive. Furthermore, are coprime, since and These imply that has a strictly primitive integral solution.
To prove the converse, suppose that an integer has a prime factor which does not split in Let be an integral solution of
Case: is unramified in or “ and is unramified in ” In the principal ideal is decomposed as
where both sides are divisible by Therefore and are divisible by in since the prime ideal divides one of these ideals and also the other because of its self-conjugacy. This implies that and are divisible by in since
(2.3) by the assumption. Hence and therefore
Case: ramifies in Then and therefore divides which implies
In these cases, is not strictly primitive. Thus the assertion of (1) holds.
Suppose that Then has no integral solutions, since
This implies that if
Suppose that Then the pair of odd integers defined by satisfies This implies that are coprime, and therefore and In this case, holds by the minimality of
Conversely, if and then a strictly primitive integral solution of satisfies
which implies since
If then
which imply that every integral solution of is a pair of even numbers for each multiple of
If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired equalities.
The desired inequality follows from the argument in the proof of (1).∎
Theorem 7.
Let be pairwise coprime ideals in the integer ring of a quadratic field Then is primitive if and only if are primitive.
Proof.
For any ideal in is primitive if and only if the residue group is cyclic (see [2, Corollary 6.30]). Combining this fact with the Chinese remainder theorem, we obtain the desired assertion. ∎
Now we are ready to prove Theorem 3.
Proof of Theorem 3.
Put and let denote the integer ring of For each prime number which splits in we take a prime ideal in lying over in and put Note that are primitive for such a prime ideal in and any Let be a strictly primitive integral solution of Let be the set of all prime factors of and put For each let and be the quotient and remainder, respectively, when dividing by For each put The equality can be expressed as with ideals in which implies
since for each For each if or then is decomposed as
in by Theorem 6 (1), and therefore the ideal is divisible by either or ; otherwise, in and also by (2.3), and therefore which contradicts that are coprime. For each if or we can see that is divisible by either or by the same argument, which implies
- (i)
Case: and Then is unramified in and by Theorem 6. The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in since
(2.4) which contradicts that are coprime. The ideal is decomposed as
with in The ideal
is principal and primitive by Theorem 7, and therefore equal to which implies that for each Hence there exists such that
Letting we obtain (1.7).
- (iii)
Case: and Then and has no strictly primitive integral solutions for each multiple of by Theorem 6, which implies The ideal is decomposed as
with in The ideal
is principal and primitive by Theorem 7, and therefore equal to which implies that for each Hence there exists such that
Letting
and defining by
where the first double sign is not necessarily the same as the second, we obtain (1.7).
The exponent satisfies (1.8), since for each and
by (see [7, Theorem 2.1.4]). Thus the assertion of (1) holds.
Let be an integral solution of and let be the greatest common divisor of and Let be the set of all prime factors of and put . Put
Then is a strictly primitive integral solution of and satisfies
(2.5) for some satisfying (1.8) and the conditions obtained by replacing with in (1.5), (1.6) because of (1). Put
Case or . Then and since
Case . Then and since and
Case or . Then
Case . Then
Multiplying both sides of (2.5) by
we obtain (1.9).∎
Remark 4.
By Theorem 3, it is impossible for both and to have strictly primitive integral solutions, if has no integral solutions.
Example 3.
The equation has the fundamental solution however, has no integral solutions. The fundamental unit of is The equations have the fundamental solutions respectively. Every integral solution of satisfies one of the following conditions for some
Theorem 3 implies the following formula for rational solutions of Pell’s equations.
Theorem 8.
Let Every rational solution of satisfies
for some such that
and
Proof.
Every integral solution of satisfies
for some
Let be a rational solution of where are coprime integers and Then is an integral solution of and satisfies
for some where
and
by Theorem 3. Dividing both sides by we obtain (2.6). Condition (2.7) follows from
and
We can remove Condition (2.8), since and for each Thus the assertion of the theorem holds. ∎
3 Rational angle bisectors
Proof of Proposition 1.
Let be a nontrivial rational solution of ( ‣ 1), and let
with Substituting these into ( ‣ 1) and multiplying both sides by we obtain
The combinations of prime factors of which divide odd times coincide with each other, since are square numbers. Let be the product of these prime factors (define if there are no such prime factors). Then there exist such that
where are the -components of the integral solutions of and are the -components of the rational solutions of with Substituting (3.2) into (3.1) and dividing both sides by we obtain
or equivalently
Solving for we obtain
since and therefore
Proof of Theorem 2.
Case : are the -components of rational solutions of respectively. Let
with Then we have
Furthermore, put
Then we have
These imply
and therefore
Case : are the -components of rational solutions of for some square-free integer respectively. In general, if and then
Combining these identities with Theorem 8, we obtain the desired assertion.∎
Appendix
In Theorem 2, for such that has a rational solution and the values of are summarized in the following table. For reference, we add the values of the ideal class number the fundamental unit of and below the values of respectively.